OpenStudy (anonymous):

what is the vertex of f(x) =x^2 + 6x -3

OpenStudy (anonymous):

vertex = -b/2a = -6/2 = -3 this is x coordinate substitute this value in f(x) to find y ordinate

OpenStudy (anonymous):

it is simple this is the equation of a parabola so the vertex is at the point where the derivative of this function is 0 so the derivative becomes 2x +6 now equate it with 0 u will get x=-3 and putting this value in the function u will get y=-12

OpenStudy (anonymous):

Thank you to both of you

OpenStudy (anonymous):

ur welcome